# Packing subgraphs in regular graphs

### v1

@@ 2.1 Balancing the expanders
Given the partition $\{Z_{1},\ldots,Z_{r}\}$ of the vertices of $G$, a natural strategy for obtaining an almost perfect $K_{t,t}$-packing in $G$ is to find an almost perfect $K_{t,t}$-packing in each of the expanders $G[Z_{i}]$. However, it is possible that some of these expanders are rather imbalanced bipartite graphs, rendering this plan impossible. To overcome this obstacle, in our first key lemma (Lemma˜4.1), we modify the sets $Z_{i}$, and remove a small collection of vertex-disjoint $K_{t,t}$-copies and a very small number of additional vertices, so that the remainder of each modified set $Z_{i}$ spans an expander which is either close to a balanced bipartite graph, or is far from being bipartite.
If the graph consisting of edges between expanders has small maximum degree, then it is relatively easy to realise this plan. Indeed, one can iteratively find vertex-disjoint $K_{t,t}$-copies in $G$ between pairs of expanders, so that the number of $K_{t,t}$’s between a given pair of expanders is proportional to the number of edges between them, using an unbalanced variant of the Kővári–Sós–Turán theorem. Removing these $K_{t,t}$-copies results in expanders that are either close enough to being balanced bipartite, or are far from being bipartite. Of course, it is possible that the maximum degree of edges between expanders is large, even after moving some vertices between expanders. Nevertheless, we emulate the small minimum degree situation as follows. We initially set $H$ to be the subgraph of edges with ends in distinct sets $Z_{i}$, $i\in[r]$, and then follow a carefully designed procedure (see Section˜4.1) that moves vertices of high degree between expanders, and modifies $H$ so as to satisfy certain invariants. After performing this procedure, $H$ ‘essentially’ has small maximum degree [^1] More precisely, the subgraph of $H$ consisting of edges between any two distinct expanders either has small maximum degree, or it has very few edges. In particular, the number of large degree vertices in $H$ is at most a constant.. Moreover, the invariants guarantee that only few vertices are moved throughout the procedure (which is essential for showing that the modified expanders are, indeed, expanders) and that $H$ still has enough edges to construct a suitable $K_{t,t}$-packing, as described (vaguely) at the beginning of this paragraph.
@@
Balancing expanders (in a different context) was also a key step in the proofs of Kühn, Lo, Osthus and Staden in [27, 28] and of Gruslys and Letzter in [11]. However, the method we develop here to construct a balanced $K_{t,t}$-packing differs significantly from these previous approaches, so it is of independent interest.
@@ 2.2 Packing $K_{t,t}$’s in expanders
Balancing the expanders (using Lemma˜4.1) reduces the original problem to the problem of finding an almost perfect $K_{t,t}$-packing (covering all but a constant number of vertices) in an expander $G$ that is either bipartite and balanced or far from bipartite. We address this reduced problem in our second key lemma, Lemma˜5.1, where we construct a perfect $K_{t,t}$-packing under the assumption that the number of vertices is divisible by $2t$, using the following approach. First, we apply the regularity lemma (Lemma˜3.5) to $G$ to obtain vertex-disjoint clusters $V_{i}$, $i\in[m]$, covering all but a small number of exceptional vertices in $G$, and a subgraph $G^{\prime}$ with vertex set $V_{1}\cup\ldots\cup V_{m}$. We will then find vertex-disjoint super-regular subgraphs $G^{\prime}[U_{2i-1},U_{2i}]$, $i\in[m]$, covering almost all vertices in $V_{1}\cup\ldots\cup V_{m}$, such that the sets $U_{1},\ldots,U_{2m}$ have the same size; for this, we reply on the Hamiltonicity of expanders to find a perfect fractional matching in the reduced graph (we explain this in more detail in the next paragraph). Next, we find three vertex-disjoint collections $\mathcal{K}_{1}$, $\mathcal{K}_{2}$, $\mathcal{K}_{3}$ of vertex-disjoint $K_{t,t}$-copies in $G$ such that $\mathcal{K}_{1}$ covers all vertices outside of $U_{1}\cup\ldots\cup U_{2m}$ while using very few vertices from each set $U_{i}$, $i\in[2m]$; $\mathcal{K}_{2}$ is very small and is chosen so that the number of vertices in $U_{i}$, $i\in[2m]$, not covered by $\mathcal{K}_{1}$ or $\mathcal{K}_{2}$ is divisible by $t$; and $\mathcal{K}_{3}$ is chosen so that the number of vertices in $U_{i}$ that are uncovered by $\mathcal{K}_{1}\cup\mathcal{K}_{2}\cup\mathcal{K}_{3}$ is the same for all $i\in[2m]$. We then apply the blow-up lemma to find a perfect $K_{t,t}$-packing in each super-regular graph $G^{\prime}[U_{2i-1}^{\prime},U_{2i}^{\prime}]$, $i\in[m]$, where $U_{i}^{\prime}$ denotes the set of uncovered vertices in $U_{i}$. These latter $K_{t,t}$-packings, together with $\mathcal{K}_{1}\cup\mathcal{K}_{2}\cup\mathcal{K}_{3}$, form the desired perfect $K_{t,t}$-packing in $G$.
For the construction of the sets $U_{1},\ldots,U_{2m}$, we consider the ‘reduced graph’ $\Gamma$ whose vertices are $V_{1},\ldots,V_{m}$, where $V_{i}$ is adjacent to $V_{j}$ if and only if $G^{\prime}[V_{i},V_{j}]$ is $\varepsilon$-regular with large enough density. We will first find a ‘perfect $2$-matching’ in $\Gamma$, that is, a collection of vertex-disjoint edges and cycles covering the vertices of $\Gamma$. This perfect $2$-matching then corresponds to a perfect matching in the ‘lift’ of $\Gamma$, so if we partition each cluster $V_{i}$ into two sets of equal size, then we obtain a perfect matching in the reduced graph whose vertices are these new clusters. Finally, by moving a small number of vertices from the new clusters into the exceptional set, we obtain the desired vertex-disjoint super-regular pairs $G^{\prime}[U_{2i-1},U_{2i}]$ for $i\in[m]$, covering all but a small fraction of the vertices in $V_{1}\cup\cdots\cup V_{m}$.
To find a perfect $2$-matching in the reduced graph $\Gamma$, it suffices to find a perfect fractional matching (see, e.g., Fact˜5.4 and its proof). Since $G^{\prime}[V_{1}\cup\cdots\cup V_{m}]$ is an almost regular expander, this in turn follows from the fact that almost regular expanders are Hamiltonian (provided they are far from bipartite, or are balanced and bipartite). The Hamiltonicity of almost regular expanders has many applications; see, for example, [27, 28, 11], and it follows from the work of Kühn, Osthus, and Treglown [34] on robust expanders. The idea of seeking perfect $2$-matchings in reduced graphs was used previously in, e.g., [17, 16, 25].
One of the main novelties in our proof of the key lemma, Lemma˜5.1, is in the construction of the collection $\mathcal{K}_{3}$ of vertex-disjoint $K_{t,t}$-copies in $G$ for ‘balancing’ the number of uncovered vertices in the sets $U_{i}$. Our strategy here is to construct a suitable matching in $G$ and use it as a template to construct the collection $\mathcal{K}_{3}$. More precisely, we consider a random subgraph $H$ (with vertex set disjoint from $V(\mathcal{K}_{1}\cup\mathcal{K}_{2})$) such that $|V(H)\cap U_{i}|=\frac{|U_{i}|}{2t}$ for every $i\in[2m]$ and show that $H$ ‘inherits’ the expansion properties of $G$. This implies that $H$ is robustly matchable, i.e. that there is a perfect matching in $H$ even after the removal of any small set of vertices (where the set we remove must balance the two parts of $H$ if it is bipartite). In particular, if $S$ is a set of vertices in $H$ such that $|S\cap U_{i}|=\frac{|V(\mathcal{K}_{1}\cup\mathcal{K}_{2})\cap U_{i}|}{t}$ for every $i\in[2m]$, then HSH-S has a perfect matching $\mathcal{M}$. We now use $\mathcal{M}$ as a template to construct the collection $\mathcal{K}_{3}$ of $K_{t,t}$’s, by picking, for every distinct $i,j\in[2m]$, as many $K_{t,t}$’s between $U_{i}$ and $U_{j}$ as there are edges of $\mathcal{M}$ between $U_{i}$ and $U_{j}$. This ensures that, for every $i\in[2m]$,
$|V(\mathcal{K}_{3})\cap U_{i}|=t\cdot|(V(H)-S)\cap U_{i}|=t\cdot\left(\frac{|U_{i}|}{2t}-\frac{|V(\mathcal{K}_{1}\cup\mathcal{K}_{2})\cap U_{i}|}{t}\right)=\frac{|U_{i}|}{2}-|V(\mathcal{K}_{1}\cup\mathcal{K}_{2})\cap U_{i}|,$
since there are $|(V(H)-S)\cap U_{i}|$ edges in $\mathcal{M}$ that are incident to the vertices of $U_{i}$. Hence, taking $U^{\prime}_{i}\coloneqq U_{i}-V(\mathcal{K}_{1}\cup\mathcal{K}_{2}\cup\mathcal{K}_{3})$, we have
$|U^{\prime}_{i}|=|U_{i}|-|V(\mathcal{K}_{1}\cup\mathcal{K}_{2})\cap U_{i}|-|V(\mathcal{K}_{3})\cap U_{i}|=\frac{|U_{i}|}{2},$
showing that the sets $U_{i}^{\prime}$ with $i\in[2m]$ all have the same size. Now, we can indeed find a perfect $K_{t,t}$-packing in each super-regular graph $G^{\prime}[U_{2i-1}^{\prime},U_{2i}^{\prime}]$ using the blow-up lemma. As discussed earlier, these $K_{t,t}$-packings together with $\mathcal{K}_{1}\cup\mathcal{K}_{2}\cup\mathcal{K}_{3}$ yield the desired perfect $K_{t,t}$-packing in $G$.
@@ 3.2 Expanders in dense regular graphs
We will use the following structural result (Lemma˜3.2) due to Gruslys and Letzter [11], which goes back to the work of Kühn, Lo, Osthus, and Staden [27, 28]. This result serves as the first step in the proofs of both Theorem˜1.4 and Theorem˜1.7. It shows that the vertices of a dense regular graph $G$ can be partitioned into a small number of sets such that the subgraphs induced by these sets exhibit strong expansion properties and have a large minimum degree. Throughout the remainder of the paper, we refer to these subgraphs as expanders. The notion of expansion used here is characterized by the absence of sparse cuts. Specifically, a graph $G$ has no $\zeta$-sparse cuts if, for every partition $\{X,Y\}$ of $V(G)$, the number of edges between $X$ and $Y$ is at least $\zeta|X||Y|$. This notion of expansion was first introduced (using different terminology and with $\zeta$ as a function of $n$) in a paper by Conlon, Fox, and Sudakov [7]. We will distinguish between expanders that are close to bipartite and those that are far from bipartite. Recall that a graph $G$ on $n$ vertices is said to be $\gamma$-almost-bipartite if one can make $G$ bipartite by removing up to $\gamma n^{2}$ edges, and otherwise, $G$ is called $\gamma$-far-from-bipartite.
@@
The following lemma is used multiple times throughout the paper. Its proof is very similar to that of Lemma 2.2 in [11], and relies on results concerning Hamiltonicity of robust expanders, due to Kühn, Osthus and Treglown [34]. The lemma shows that expanders far from being bipartite are Hamiltonian and remain Hamiltonian even after the removal of any small set of vertices. However, expanders that are bipartite may not be Hamiltonian – for example, an imbalanced bipartite graph cannot be Hamiltonian. In this case, the lemma shows that such expanders still become Hamiltonian after the removal of any small set of vertices that balances its two parts. We prove Lemma˜3.3 in Appendix˜B.
@@
We use the following lemma from [11] in the proof of Theorem 1.7. When presented with a partition of a dense regular graph into expanders (by applying Lemma 3.2), Lemma˜3.4 produces a small linear forest (i.e., a small collection of vertex-disjoint paths) whose removal balances the expanders that are close to being bipartite. A similar linear forest was also constructed in [27, 28].

### v2

@@ 2.1 Balancing the expanders
Given the partition $\{Z_{1},\ldots,Z_{r}\}$ of the vertices of $G$, a natural strategy for obtaining an almost perfect $K_{t,t}$-packing in $G$ is to find an almost perfect $K_{t,t}$-packing in each of the expanders $G[Z_{i}]$. However, it is possible that some of these expanders are rather imbalanced bipartite graphs, rendering this plan impossible. To overcome this obstacle, in our first key lemma (Lemma 4.1), we modify the sets $Z_{i}$ and remove a small collection of vertex-disjoint $K_{t,t}$-copies and a very small number of additional vertices so that the remainder of each modified set $Z_{i}$ spans an expander that is either bipartite and balanced or is far from bipartite. (In other words, this ‘balances’ the expanders $G[Z_{i}]$ that are almost bipartite.)
The first idea towards balancing the almost bipartite expanders is to use the regularity of $G$. For the rest of this section, let us assume for simplicity that $G$ itself is bipartite, and denote the bipartition of $G$ by $\{X,Y\}$ and the corresponding bipartition of $Z_{i}$ by $\{X_{i},Y_{i}\}$. Write $V:=V(G)$, and let $H$ denote the subgraph of $G$ on vertex set $V$ consisting of all edges whose endpoints lie in distinct sets $Z_{i}$.
A simple counting argument, using that $G$ is $d$-regular, shows that the following holds for every $i\in[r]$.
$e_{H}(X_{i},V-Z_{i})-e_{H}(Y_{i},V-Z_{i})=e_{G}(X_{i},V-Z_{i})-e_{G}(Y_{i},V-Z_{i})=(|X_{i}|-|Y_{i}|)\cdot d.$ (1)
Pretending for a moment that $t=1$, so that a $K_{t,t}$-packing is just a matching, a natural aim is thus to find a matching consisting of $\frac{1}{d}e_{H}(X_{i},V-Z_{i})$ edges leaving $Z_{i}$ and incident to $X_{i}$, and $\frac{1}{d}e_{H}(Y_{i},V-Z_{i})$ edges leaving $Z_{i}$ and incident to $Y_{i}$. By (1), this guarantees that the matching leaves the same number of vertices uncovered in $X_{i}$ and $Y_{i}$. [^1] We ignore rounding errors due to taking ceilings or floors; this is justified as we are allowed $O(1)$ uncovered vertices. To achieve this simultaneously for all $i\in[r]$, it suffices to find a matching consisting of $\frac{1}{d}e_{H}(X_{i},Y_{j})$ edges between $X_{i}$ and $Y_{j}$, for every distinct $i,j\in[r]$.
It turns out that it is quite easy to construct such a matching if all bipartite subgraphs $H[X_{i},Y_{j}]$ for $i\neq j$ have maximum degree at most $\rho n$, for some small constant $0<\rho\ll c$. Indeed, we consider the pairs $(X_{i},Y_{j})$ with $i\neq j$ one by one, in increasing order of $e_{H}(X_{i},Y_{j})$, and at each step we construct a matching of size $\frac{1}{d}e_{H}(X_{i},Y_{j})$ in $H[X_{i},Y_{j}]$ that is vertex-disjoint from all previously defined matchings. We claim that this is always possible. Indeed, consider a pair $(X_{i},Y_{j})$. Denote by $e$ the number of edges between $X_{i}$ and $Y_{j}$ and by $X_{i}^{\prime}\subseteq X_{i}$ and $Y_{j}^{\prime}\subseteq Y_{j}$ the sets of vertices that were used in previously defined matchings. Then $|X_{i}^{\prime}|,|Y_{j}^{\prime}|\leq r^{2}\cdot e/d$, by the order in which we process the pairs. By the maximum degree assumption and since $d\geq cn$, we have
$e_{H}(X_{i}-X_{i}^{\prime},Y_{j}-Y_{j}^{\prime})\geq e_{H}(X_{i},Y_{j})-\rho n\cdot(|X_{i}^{\prime}|+|Y_{j}^{\prime}|)\geq e-e/2=e/2.$
Since every graph $H^{\prime}$ has a matching of size at least $e(H^{\prime})/2\Delta(H^{\prime})$, it follows that there is a matching of size at least $e/(4\rho n)\geq e/d$ in $H[X_{i}-X_{i}^{\prime},,Y_{j}-Y_{j}^{\prime}]$, as required.
For general $t\geq 1$, our aim is the same, except that we replace matching edges between $X_{i}$ and $Y_{j}$ with copies of $K_{t,t}$ in $G$, each consisting of one vertex in $X_{i}$, $t$ vertices in $Y_{j}$, and t1t-1 vertices in $X_{j}$, so as to achieve the same balancing effect as moving one vertex from $X_{i}$ to $X_{j}$. The above greedy algorithm still applies if we use an asymmetric version of the Kővári–Sós–Turán theorem (Lemma 3.8) for all pairs $(X_{i},Y_{j})$ such that $H[X_{i},Y_{j}]$ contains at least $Cn$ edges, where $C$ is a large constant. Since we allow $O(1)$ uncovered vertices, we may ignore pairs $(X_{i},Y_{j})$ with fewer edges. Indeed, this yields a ‘balancing’ $K_{t,t}$-packing $\mathcal{K}$, which satisfies $|X_{i}-V(\mathcal{K})|=|Y_{i}-V(\mathcal{K})|+O(1)$ for all $i$.
The main challenge in proving Lemma 4.1 is therefore to emulate the situation in which the maximum degree (between the sets $Z_{i}$) is bounded by $\rho n$, which we achieve as follows. We perform the procedure below, sequentially modifying the graph $H$ and the sets $Z_{1},\ldots,Z_{r}$. If there is a pair $(X_{i},Y_{j})$ with $e_{H}(X_{i},Y_{j})\geq d$ and a vertex $v\in X_{i}$ that has at least $\rho n$ neighbours in $H$ lying in $Y_{j}$, then we perform the following three operations: we move $v$ from $X_{i}$ to $X_{j}$; we remove all edges between $v$ and $Y_{j}$ from $H$; and we remove $d-d_{H}(v)$ edges between $X_{i}-{v}$ and $Y_{j}$ from $H$, chosen arbitrarily. (The fact that the third operation is possible follows from the assumption that $e_{H}(X_{i},Y_{j})\geq d$.) Note that this procedure only removes edges from $H$ in each step and never adds any. In particular, although $v$ is now moved to $X_{j}$ we do not add to $H$ the edges from $v$ to $Y_{i}$. This shows that the number of edges in $H$ decreases by at least $\rho n$ in each step of the procedure. The first two operations are quite natural: since $v$ has large degree into $Y_{j}$, adding it to $X_{j}$ would preserve the expansion properties of $G[X_{j},Y_{j}]$. Since $H$ initially contained no edges between $X_{\ell}$ and $Y_{\ell}$ for any $\ell$, the second operation preserves this condition after each step of the procedure. The motivation for the third operation is the observation that if $H,X_{1},\ldots,X_{r},Y_{1},\ldots,Y_{r}$ satisfy (1) and we modify them by performing the three operations above, then the resulting $H,X_{1},\ldots,X_{r},Y_{1},\ldots,Y_{r}$ still satisfy (1). Indeed, for example, to see that (1) continues to hold for $X_{i},Y_{i}$, note that after performing the three operations above, both the left-hand and right-hand sides of (1) decrease by exactly $d$.
Continuing this procedure for as long as possible (also with the roles of $X_{i}$ and $Y_{i}$ reversed), the resulting graph $H$ and sets $X_{1},\ldots,X_{r},Y_{1},\ldots,Y_{r}$ still satisfy (1), and for every $i\neq j$, the bipartite subgraph $H[X_{i},Y_{j}]$ either has maximum degree at most $\rho n$ or contains fewer than $d$ edges. Therefore, this allows us to run the above procedure on $H$ to find a $K_{t,t}$-packing $\mathcal{K}$ that balances the modified sets $Z_{i}$ so that $|X_{i}-V(\mathcal{K})|=|Y_{i}-V(\mathcal{K})|+O(1)$ for every $i$. Moreover, since $H$ initially has few edges (as there are few edges between the sets $Z_{i}$) and each step removes at least $\rho n$ edges, we are guaranteed that only few vertices are moved throughout the procedure; this is essential for showing that the modified sets $Z_{i}$ are still expanders.
@@
Balancing expanders (in different contexts) was also a key step in the proofs of Kühn, Lo, Osthus and Staden in [27, 28] and of Gruslys and Letzter in [11]. The method we develop here to construct a balanced $K_{t,t}$-packing differs significantly from these previous approaches.
@@ 2.2 Packing $K_{t,t}$’s in expanders
Having balanced the expanders (using Lemma 4.1), the original problem is reduced to the problem of finding an almost perfect $K_{t,t}$-packing (covering all but a constant number of vertices) in an expander $G$ that is dense, almost regular, and either far from bipartite or balanced and bipartite. We address this reduced problem in our second key lemma, Lemma 5.1, where we construct a perfect $K_{t,t}$-packing in such an expander under the assumption that the number of vertices is divisible by $2t$.
The proof of Lemma 5.1 uses the regularity lemma and combines two main themes. First, we exploit the fact that dense, almost regular expanders that are either far from bipartite or bipartite and balanced contain Hamilton cycles, following the work of Kühn, Osthus, and Treglown [34] on robust expanders. Second, as in the previous section, we again use $K_{t,t}$-packings for balancing purposes, but now in two distinct ways that differ from the approach used earlier.
To begin, we aim to find disjoint sets $U_{1},\ldots,U_{2m}$ of equal size that cover almost all vertices of $G$, such that almost every pair $(U_{i},U_{j})$ is $\varepsilon$-regular in $G$, and each pair $G[U_{2i-1},U_{2i}]$ is somewhat dense and super-regular (that is, almost every pair $(U_{i},U_{j})$ is $\varepsilon$-regular in $G$, and moreover $G[U_{2i-1},U_{2i}]$ has reasonably large minimum degree). To achieve this, we apply the regularity lemma (Lemma 3.5) to $G$ to obtain vertex-disjoint clusters $V_{1},\ldots,V_{m}$ of equal size that cover almost all vertices of $G$ and such that almost all pairs are $\varepsilon$-regular. We then use the expansion properties of $G$, together with the aforementioned result on Hamiltonicity of expanders, to deduce that the reduced graph $\Gamma$ (whose vertices are $V_{1},\ldots,V_{m}$ and whose edges correspond to dense $\varepsilon$-regular pairs) admits a perfect fractional matching. This, in turn, implies the existence of a perfect $2$-matching in $\Gamma$, that is, a collection of vertex-disjoint edges and cycles covering all vertices. Consequently, splitting each cluster $V_{i}$ into two equal parts yields a perfect matching in the corresponding reduced graph. By possibly removing a small proportion of vertices from each of these new clusters, we can ensure that each matched pair is super-regular, as required. The idea of seeking perfect $2$-matchings in reduced graphs has appeared previously, for example in [17, 16, 25].
Having found the sets $U_{1},\ldots,U_{2m}$, we can apply the blow-up lemma to each pair $G[U_{2i-1},U_{2i}]$ to cover all but $2t$ vertices of $U_{2i-1}\cup U_{2i}$ with a $K_{t,t}$-packing. However, this approach may only produce a $K_{t,t}$-packing in $G$ that leaves linearly many vertices uncovered, namely those outside the sets $U_{1},\ldots,U_{2m}$. To cover all vertices of $G$ (rather than almost all), we instead find a $K_{t,t}$-packing $\mathcal{K}$ that covers every vertex outside $U_{1}\cup\cdots\cup U_{2m}$ and only very few vertices in each $U_{i}$, and, crucially, leaves the same number of uncovered vertices in each $U_{i}$, where this number is divisible by $t$. Once such a $K_{t,t}$-packing $\mathcal{K}$ is found, we can apply the blow-up lemma to cover the remaining vertices in each pair $G[U_{2i-1},U_{2i}]$ with a perfect $K_{t,t}$-packing, thereby completing $\mathcal{K}$ to a perfect $K_{t,t}$-packing in $G$.
To construct such a $K_{t,t}$-packing $\mathcal{K}$, we proceed in three steps. First, we construct a packing $\mathcal{K}_{1}$ that covers all vertices outside $U_{1}\cup\cdots\cup U_{2m}$ while using very few vertices from each set $U_{i}$. This is straightforward using the Kővári–Sós–Turán theorem together with the properties of $G$. Next, we construct two further $K_{t,t}$-packings, $\mathcal{K}_{2}$ and $\mathcal{K}_{3}$, each serving a different balancing role. The packing $\mathcal{K}_{2}$ is chosen so that, for every $i$, the number of vertices in $U_{i}$ uncovered by $\mathcal{K}_{1}\cup\mathcal{K}_{2}$ is divisible by $t$. We then construct $\mathcal{K}_{3}$ so that the number of vertices in each $U_{i}$ uncovered by $\mathcal{K}_{1}\cup\mathcal{K}_{2}\cup\mathcal{K}_{3}$ is the same for all $i$, and still divisible by $t$. Taking $\mathcal{K}:=\mathcal{K}_{1}\cup\mathcal{K}_{2}\cup\mathcal{K}_{3}$ allows us to complete the proof using the blow-up lemma, as explained in the previous paragraph.
For the construction of $\mathcal{K}_{2}$, the key idea is that a walk $U_{i_{1}}\ldots U_{i_{2s-1}}$ of even length in the reduced graph allows us to “transfer” vertices from $U_{i_{1}}$ to $U_{i_{2s-1}}$ (modulo $t$). More precisely, we construct a collection of copies of $K_{t,t}$ between the clusters corresponding to this walk so as to create a balancing effect on the clusters at the two ends of the walk when these copies are removed. By combining several such walks, we can construct a collection $\mathcal{K}_{2}$ of copies of $K_{t,t}$ such that the number of uncovered vertices in each $U_{i}$ is divisible by $t$, as desired (see Section 5.4 for further details).
We now describe the construction of $\mathcal{K}_{3}$. For each $i\in[2m]$, let $U_{i}^{\prime}$ denote the set of vertices in $U_{i}$ that remain uncovered. Our first step is to find a template matching $\mathcal{M}$ in $G[U_{1}^{\prime}\cup\ldots\cup U_{2m}^{\prime}]$ such that $|U_{i}^{\prime}|-t\cdot|V(\mathcal{M})\cap U_{i}^{\prime}|$ is the same for all $i\in[2m]$. Given such a matching $\mathcal{M}$, we apply the blow-up lemma to construct a $K_{t,t}$-packing $\mathcal{K}_{3}$ by replacing each edge of $\mathcal{M}$ between $U_{i}^{\prime}$ and $U_{j}^{\prime}$ with a copy of $K_{t,t}$ in $G[U_{i}^{\prime},U_{j}^{\prime}]$. Consequently, the number of vertices in $U_{i}^{\prime}$ that remain uncovered by $\mathcal{K}_{3}$ is exactly $|U_{i}^{\prime}|-t\cdot|V(\mathcal{M})\cap U_{i}^{\prime}|$. By the choice of $\mathcal{M}$, this quantity is the same for all $i\in[2m]$; moreover, it is divisible by $t$ by choice of $\mathcal{K}_{2}$, as required. Finally, to construct the template matching $\mathcal{M}$, we select subsets $U_{i}^{\prime\prime}\subseteq U_{i}$ of appropriate size and show that the induced graph $G[U_{1}^{\prime\prime}\cup\ldots\cup U_{2m}^{\prime\prime}]$ inherits the relevant expansion properties of $G$. By the aforementioned Hamiltonicity of expanders, this graph contains a perfect matching, which serves as the desired template matching $\mathcal{M}$ (see Section 5.5 for further details).
@@ 3.2 Expanders in dense regular graphs
We will use the following structural result (Lemma 3.2) due to Gruslys and Letzter [11], which goes back to the work of Kühn, Lo, Osthus, and Staden [27, 28]. This result serves as the first step in the proofs of both Theorem 1.4 and Theorem 1.7. It shows that the vertices of a dense regular graph $G$ can be partitioned into a small number of sets such that the subgraphs induced by these sets exhibit strong expansion properties and have a large minimum degree. Throughout the remainder of the paper, we refer to these subgraphs as expanders. The notion of expansion used here is characterized by the absence of sparse cuts. Specifically, a graph $G$ has no $\zeta$-sparse cuts if, for every partition $\{X,Y\}$ of $V(G)$, the number of edges between $X$ and $Y$ is at least $\zeta|X||Y|$. This notion of expansion was first introduced (using different terminology and with $\zeta$ as a function of $n$) in a paper by Conlon, Fox, and Sudakov [7]. We will distinguish between expanders that are close to bipartite and those that are far from bipartite. Recall that a graph $G$ on $n$ vertices is said to be $\gamma$-almost-bipartite if one can make $G$ bipartite by removing up to $\gamma n^{2}$ edges, and otherwise, $G$ is called $\gamma$-far-from-bipartite.
@@
The following lemma is used multiple times throughout the paper. Its proof is very similar to that of Lemma 2.2 in [11], and relies on results concerning Hamiltonicity of robust expanders, due to Kühn, Osthus and Treglown [34]. The lemma shows that expanders far from being bipartite are Hamiltonian and remain Hamiltonian even after the removal of any small set of vertices. However, expanders that are bipartite may not be Hamiltonian – for example, an imbalanced bipartite graph cannot be Hamiltonian. In this case, the lemma shows that such expanders still become Hamiltonian after the removal of any small set of vertices that balances its two parts. We prove Lemma 3.3 in Appendix B.
@@
We use the following lemma from [11] in the proof of Theorem 1.7. When presented with a partition of a dense regular graph into expanders (by applying Lemma 3.2), Lemma 3.4 produces a small linear forest (i.e., a small collection of vertex-disjoint paths) whose removal balances the expanders that are close to being bipartite. A similar linear forest was also constructed in [27, 28].